### v4

@@ Abstract
In recent decades, several assistive technologies have been developed to improve the ability of blind and visually impaired individuals to navigate independently and safely. At the same time, simultaneous localization and mapping (SLAM) techniques have become sufficiently robust and efficient to be adopted in developing these assistive technologies. We present the first systematic literature review of 54 recent studies on SLAM-based solutions for blind and visually impaired people, focusing on literature published from 2017 onward. This review explores various localization and mapping techniques employed in this context. We systematically identified and categorized diverse SLAM approaches and analyzed their localization and mapping techniques, sensor types, computing resources, and machine-learning methods. We discuss the advantages and limitations of these techniques for blind and visually impaired navigation. Moreover, we examine the major challenges described across studies, including practical considerations that affect usability and adoption. Our analysis also evaluates the effectiveness of these SLAM-based solutions in real-world scenarios and user satisfaction, providing insights into their practical impact on BVI mobility. The insights derived from this review identify critical gaps and opportunities for future research activities, particularly in addressing the challenges presented by dynamic and complex environments. We explain how SLAM technology offers the potential to improve the ability of visually impaired individuals to navigate effectively. Finally, we present future opportunities and challenges in this domain.
@@ I Introduction
The collaborative effort between the front- and back-ends empowers SLAM to provide a robust and real-time spatial understanding, making it a valuable tool for various applications. The SLAM community has made tremendous strides over the past 35 years, developing large-scale practical applications and seeing a steady transition of this technology into the industry [3]. Cadena et al. [3] divided the life cycle of SLAM into three periods: classical age (1986-2004), algorithmic analysis age (2004–2015), and robust-perception age (2015-present). The evolution of portable computation and the availability of low-cost, highly accurate, and lightweight sensors such as cameras and IMUs make them appropriate for pedestrian navigation. By exploiting these advances, many researchers have recently adopted SLAM to develop assistive technology demonstrators to help BVI people navigate unknown environments.
@@ III Result
Since there are many different types of SLAMs designed with different types of sensors, for different applications, and for different scenarios, in this section, we focus on reviewing the types of SLAM that have been used for the navigation of visually impaired people. It is a key technology in robotics and computer vision and has the potential to assist visually impaired individuals with navigation. This can help visually impaired individuals provide real-time location information, maps, and spatial awareness. Among the 54 studies surveyed, three strategies were common, as shown in Table XVI. In this table, we use the exact terms mentioned in the literature for the localization and mapping techniques.
@@ III-C2 Practical challenges and considerations
In addition to the technical and methodological aspects, we recognized the importance of practical challenges considerations that can significantly affect the usability and adoption of SLAM-based assistive systems. Therefore, we included an evaluation of the practical challenges and operational efficiency, as summarized in Tables XXIV and XXV. The information in these tables has either been directly extracted from the article or can be easily inferred from the article’s text. These tables provide information on user-friendliness, cost-efficiency, weight, comfort for extended use, adjustable fit, fatigue mitigation, and portability of the assistive tools described in the reviewed studies. For instance, while smartphones and lightweight devices such as eyeglasses-mounted sensors [74] and ARCore-supported smartphones with haptic gloves [75] are generally well received because of their high portability and ease of use, heavier devices such as guiding robots [68] and rolling suitcase-shaped device [95] are noted to cause user fatigue over extended periods. The augmented cane [67], although found to improve confidence and workload for novice and expert users, also faced usability challenges owing to its weight. Clear instructions and easy learning curves, as seen in electronic glasses with haptic modules [57], play a significant role in enhancing user satisfaction. However, the cost efficiency of these technologies varies, with some solutions being more affordable and accessible.
@@ III-D2 Effectiveness of localization and mapping techniques in real-world scenarios
Many studies, such as [57], [59], and [62], have demonstrated high localization and mapping accuracy, particularly in indoor environments. These studies employed techniques such as ORB-SLAM2, OpenVSLAM, and Cartographer to ensure reliable feature matching and adaptive navigation.
@@
[82] High: Global pose estimation within 80 cm radius of ground truth; verified in real-world scenarios. Low: Navigation relies on SLAM-relative poses, prone to instability with changing or occluded features. $\bullet$ Real-time
@@ IV Future opportunities
Challenge scenarios and real-world studies Navigating crowded environments remains a significant challenge for the visually impaired and studies addressing this issue are limited. Evaluations are often conducted in controlled settings rather than real-world scenarios. Future research should focus on developing and testing solutions in high-traffic public places such as train stations and shopping malls. Furthermore, addressing challenging conditions, such as changes in illumination, low-light scenarios, high-speed dynamic objects, and complex backgrounds, can enhance the robustness and versatility of navigation systems. Techniques such as image enhancement for ORB points and LSD line feature recovery used in agricultural environments [116] can be adapted for visually impaired navigation.
Future research should prioritize real-world testing and validation by conducting extensive field tests in high-traffic public areas, such as train stations, airports, and shopping centers, to evaluate the system performance under diverse environmental factors. Handling dynamic and complex environments requires the development of algorithms that can adapt to rapid changes such as sudden variations in lighting, moving objects, and different background textures. Machine learning models that can learn and predict common patterns in dynamic environments can improve system responsiveness and reliability.
To enhance robustness under low-light and variable lighting conditions, one solution is to integrate advanced image enhancement techniques and utilize sensors complementary to traditional cameras, ensuring better obstacle detection and navigation assistance in poorly lit areas. It is essential that these enhancements do not compromise the real-time performance of the system and maintain seamless and immediate feedback for the user.
@@
Long-term navigation The development of solutions that are effective over extended navigation periods is critical to achieve autonomous navigation. These solutions must ensure accurate mapping and localization even when the maps are updated over a longer navigation duration. To address this challenge, researchers can leverage the proposed solutions in robotics, such as that presented in [117], which introduced a novel long-term SLAM system with map prediction and dynamic removal, thereby allowing wheelchair robots to maintain precise navigation capabilities over extended periods. Future research should focus on developing robust algorithms for continuous map updates and maintenance, including strategies for handling environmental changes over time. This involves dynamic object removal and adaptive map refinement to ensure the accuracy and relevance of navigation maps.
@@
Deep learning integration The integration of deep learning with SLAM algorithms for BVI navigation requires further investigation. Deep learning offers a versatile approach for enhancing various aspects of SLAM such as precise pose estimation under challenging conditions, relocalization, and loop-closure detection. Despite challenges, such as the need for large, accurately labeled datasets, the black-box nature of deep-learning models, and the computational intensity of SLAM systems, the association between deep learning and SLAM holds promise for advancing navigation solutions for the visually impaired, particularly in challenging scenarios. Future research should focus on developing more efficient deep-learning models that can operate effectively with limited computational resources and real-time constraints. In addition, creating large-scale, accurately labeled datasets tailored for BVI navigation is crucial for training robust models. Addressing the interpretability of deep learning models can also enhance the trust and transparency in these systems. Collaboration among machine-learning experts, roboticists, and vision scientists can drive the development of innovative algorithms that leverage deep learning to enhance the reliability and accuracy of SLAM-based navigation aids for the visually impaired.
Indoor and outdoor navigation integration Seamless transitions between indoor and outdoor environments are crucial for enhancing the independence and mobility of BVI individuals. However, most studies in our SLR have focused primarily on indoor environments. Future research should aim to develop solutions that provide unified and consistent navigation experience in both indoor and outdoor settings. Researchers should explore the integration of robust sensor fusion techniques and adaptive algorithms capable of handling the different conditions and challenges of these environments. Additionally, extensive real-world testing in various settings is essential to ensure the practicality and reliability of these solutions.
@@
Obstacle detection Achieving detailed knowledge of obstacles and their characteristics is essential for BVI people. Although some studies included in the SLR addressed obstacle detection, the depth and accuracy of the obstacle information provided may still be limited. To address the need for more detailed and accurate obstacle information for BVI individuals, future research should focus on advancing SLAM algorithms to deliver context-aware obstacle detection. This involves integrating semantic understanding with precise spatial measurements, allowing the system to identify and interpret the nature and significance of obstacles accurately.
Researchers should explore methods to enhance the real-time processing capabilities to ensure that the system can operate effectively in dynamic and complex environments. Additionally, it is crucial to develop algorithms that can learn and adapt to various obstacle types and scenarios through continuous data collection and machine learning.
@@
Drawing inspiration from approaches used in robotics and autonomous drones, such as the real-time metric-semantic SLAM demonstrated by [118], can provide valuable insights. This approach integrates semantic understanding with precise distance and spatial relationship measurements, enabling more comprehensive and reliable navigation aids. Therefore, future research should prioritize improving both the depth and accuracy of obstacle information, while ensuring robust real-time performance and adaptability to various real-world conditions.
@@
Semantic information integration Integrating semantic information into SLAM algorithms can significantly enhance the performance and robustness of navigation systems for BVI individuals. This information can be used to refine the mapping and localization processes and enhance the overall reliability of navigation in complex environments. For instance, semantic information can help reject outliers in loop closure detection, which is a crucial step in SLAM, which involves identifying and matching previously visited locations.
To advance this area, future research should focus on developing advanced techniques for semantic data extraction and integration within the SLAM frameworks. Researchers should also explore methods to ensure real-time performance while maintaining the accuracy and detail of the semantic information.
@@
Product development and collaboration Notably, all the reviewed approaches were prototypes in the early stages of research and are not yet practical. This might be due to the absence of a unified community or group dedicated to solving the BVI navigation challenges. Much of the work in this domain has been carried out by academic groups or small companies that often fail to produce feasible final products. This underscores a significant future opportunity to develop collaboration and to bridge the gap between research and practical implementation. Additionally, efforts should be made to develop standardized evaluation metrics and protocols to ensure that the developed systems meet real-world needs and can be effectively transitioned from prototypes to market-ready solutions. Encouraging partnerships with technology companies can also accelerate the commercialization process. These partnerships provide the necessary support to bring innovative solutions to the market.
@@ V Conclusion
We investigated the challenging scenarios encountered by SLAM-based navigation systems, which have been addressed in the literature. Additionally, we discussed practical considerations that affect the usability and adoption of these systems. Furthermore, we analyzed how the proposed SLAM-based solutions improve the mobility and navigation of visually impaired individuals. We evaluated the effectiveness of these solutions in real-world scenarios and assessed the user satisfaction to understand their practical impact on BVI mobility. Finally, we identified gaps, opportunities, and areas of interest that could be explored further in future research, such as addressing challenges in crowded environments, improving real-world applicability, integrating deep learning, and ensuring long-term navigation effectiveness in SLAM-based solutions for visually impaired navigations.

### v5

@@ Abstract
In recent decades, several assistive technologies have been developed to improve the ability of blind and visually impaired individuals to navigate independently and safely. At the same time, simultaneous localization and mapping (SLAM) techniques have become sufficiently robust and efficient to be adopted in developing these assistive technologies. We present the first systematic literature review of 54 recent studies on SLAM-based solutions for blind and visually impaired people, focusing on literature published from 2017 onward. This review explores various localization and mapping techniques employed in this context. We systematically identified and categorized diverse SLAM approaches and analyzed their localization and mapping techniques, sensor types, computing resources, and machine-learning methods. We discuss the advantages and limitations of these techniques for blind and visually impaired navigation. Moreover, we examine the major challenges described across studies, including practical challenges and considerations that affect usability and adoption. Our analysis also evaluates the effectiveness of these SLAM-based solutions in real-world scenarios and user satisfaction, providing insights into their practical impact on BVI mobility. The insights derived from this review identify critical gaps and opportunities for future research activities, particularly in addressing the challenges presented by dynamic and complex environments. We explain how SLAM technology offers the potential to improve the ability of visually impaired individuals to navigate effectively. Finally, we present future opportunities and challenges in this domain.
@@ I Introduction
The collaborative effort between the front- and back-ends empowers SLAM to provide a robust and real-time spatial understanding, making it a valuable tool for various applications. The SLAM community has made tremendous strides over the past 35 years, developing large-scale practical applications and seeing a steady transition of this technology into the industry [3].
SLAM technology has been widely applied in various fields, demonstrating its versatility and robustness. In autonomous driving, SLAM enables vehicles to create and update maps of their environments in real time [130]. This capability is essential for safe and efficient road navigation, allowing accurate localization and obstacle avoidance in dynamic urban environments. In robotics, SLAM is crucial for navigation [131]. This allows robots to operate effectively in unknown environments by simultaneously building a map of their surroundings and determining their positions within them. Such a capability is essential for autonomous exploration, path planning, and obstacle avoidance in diverse settings, ranging from indoor spaces to outdoor terrains. Augmented reality (AR) relies heavily on SLAM for the accurate placement of virtual objects in the real world, enhancing user experiences in gaming, education, and industrial applications [132]. In underwater robotics, SLAM helps autonomous underwater vehicles navigate highly unstructured and complex marine environments and supports exploration, research, and maintenance [133]. For aerial vehicles, SLAM is an indispensable methodology for autonomous flight performed by unmanned aerial vehicles (UAVs) [134], along with flight control [135, 136]. SLAM enables drones to navigate and map areas autonomously, which is valuable for tasks such as search-and-rescue, surveying, and environmental monitoring. These diverse applications highlight the versatility of SLAM and its critical role in enabling autonomous operation across various scenarios and platforms, from urban landscapes to ocean depths.
The evolution of portable computation and the availability of low-cost, highly accurate, and lightweight sensors such as cameras and IMUs make them appropriate for pedestrian navigation. By exploiting these advances, many researchers have recently adopted SLAM to develop assistive technology demonstrators to help BVI people navigate unknown environments.
@@ III Result
Given the variety of SLAM systems designed with different sensors, applications, and scenarios, this section focuses specifically on reviewing the types of SLAM used for the navigation of the visually impaired. It is a key technology in robotics and computer vision and has the potential to assist visually impaired individuals with navigation. This can help visually impaired individuals provide real-time location information, maps, and spatial awareness. Among the 54 studies surveyed, three strategies were common, as shown in Table XVI. In this table, we use the exact terms mentioned in the literature for the localization and mapping techniques.
@@ III-C2 Practical challenges and considerations
In addition to the technical and methodological aspects, we recognized the importance of practical challenges and considerations that can significantly affect the usability and adoption of SLAM-based assistive systems. Therefore, we included an evaluation of the practical challenges and operational efficiency, as summarized in Tables XXIV and XXV. The information in these tables has either been directly extracted from the article or can be easily inferred from the article’s text. These tables provide information on user-friendliness, cost-efficiency, weight, comfort for extended use, adjustable fit, fatigue mitigation, and portability of the assistive tools described in the reviewed studies. For instance, while smartphones and lightweight devices such as eyeglasses-mounted sensors [74] and ARCore-supported smartphones with haptic gloves [75] are generally well received because of their high portability and ease of use, heavier devices such as guiding robots [68] and rolling suitcase-shaped device [95] are noted to cause user fatigue over extended periods. The augmented cane [67], although found to improve confidence and workload for novice and expert users, also faced usability challenges owing to its weight. Clear instructions and easy learning curves, as seen in electronic glasses with haptic modules [57], play a significant role in enhancing user satisfaction. However, the cost efficiency of these technologies varies, with some solutions being more affordable and accessible.
@@ III-D2 Effectiveness of localization and mapping techniques in real-world scenarios
Many studies, such as [57], [59], and [62], have demonstrated high localization and mapping accuracy, particularly in indoor environments. These studies employed techniques such as ORB-SLAM2 and Cartographer to ensure reliable feature matching and adaptive navigation.
@@
[82] Moderate: Global pose estimation within 80 cm radius of ground truth; verified in real-world scenarios. Low: Navigation relies on SLAM-relative poses, prone to instability with changing or occluded features. $\bullet$ Real-time
@@ IV Future opportunities
Challenge scenarios and real-world studies Navigating crowded environments remains a significant challenge for the visually impaired and studies addressing this issue are limited. Evaluations are often conducted in controlled settings rather than real-world scenarios. This is problematic because controlled environments may not accurately reflect the dynamic and unpredictable nature of real-world crowded spaces, where visually impaired individuals face numerous obstacles and safety risks. Future research should focus on developing and testing solutions in high-traffic public places such as train stations and shopping malls. This is crucial to ensure that assistive navigation systems can effectively handle the complexities of real-world crowded environments, including the presence of numerous dynamic objects, varying crowd densities, and unpredictable pedestrian behaviors.
Furthermore, addressing challenging conditions such as changes in illumination, low-light scenarios, high-speed dynamic objects, and complex backgrounds can enhance the robustness and versatility of navigation systems. These challenging conditions are common in real-world scenarios and can significantly impact the performance and reliability of SLAM-based navigation systems. By addressing these challenges, researchers can develop more robust and adaptable solutions that can function effectively in diverse and demanding environments.
Techniques such as image enhancement for ORB points and LSD line feature recovery used in agricultural environments [116] can be adapted for visually impaired navigation. Adapting these techniques from other domains can accelerate the development of more effective solutions for visually impaired navigation, leveraging existing knowledge and expertise to address the specific challenges faced by this user group.
Long-term navigation The development of solutions that are effective over extended navigation periods is critical to achieve autonomous navigation. These solutions must ensure accurate mapping and localization even when the maps are updated over a longer navigation duration. This is important because environments are not static; they change over time. Obstacles may appear or disappear, and landmarks may be altered. A SLAM-based navigation system that cannot adapt to these changes will become increasingly inaccurate and unreliable over time, potentially leading to dangerous situations for visually impaired users.
@@
To address this challenge, researchers can leverage solutions proposed in robotics, such as those presented in [117], which introduced a novel long-term SLAM system with map prediction and dynamic removal, thereby allowing wheelchair robots to maintain precise navigation capabilities over extended periods.
Future research should focus on the development of robust algorithms for continuous map updates and maintenance, including strategies for handling environmental changes over time. These strategies are essential to ensure that the navigation system can maintain its accuracy and reliability over extended periods, even in the face of environmental changes. By continuously updating and refining the map, the system can provide visually impaired users with up-to-date and relevant information about their surroundings, enabling them to navigate safely and confidently.
@@
Deep learning integration The integration of deep learning with the SLAM algorithms for BVI navigation requires further investigation. Deep learning offers a versatile approach for enhancing various aspects of SLAM such as precise pose estimation under challenging conditions, relocalization, and loop-closure detection. This is important because deep learning can potentially improve the accuracy and robustness of SLAM in complex and dynamic environments, where traditional SLAM algorithms may struggle. For instance, deep learning can be used to improve feature detection and matching in low-light conditions or to predict and adapt to changes in the environment. Despite challenges, such as the need for large, accurately labeled datasets, the black-box nature of deep-learning models, and the computational intensity, the association between deep learning and SLAM holds promise for advancing navigation solutions for the visually impaired, particularly in challenging scenarios. The potential benefits of deep learning for SLAM are substantial, and overcoming these challenges could lead to significant advancements in assistive navigation technology.
Future research should focus on developing more efficient deep learning models that can operate effectively with limited computational resources and real-time constraints. This is crucial because visually impaired individuals need real-time feedback and guidance to navigate safely and effectively. Deep learning models that are computationally intensive or require powerful hardware may not be practical for real-world use. Additionally, creating large-scale, accurately labeled datasets tailored for BVI navigation is crucial for training robust models. The lack of such datasets is a major obstacle to the development of effective deep-learning-based SLAM systems for visually impaired navigation.
Addressing the interpretability of deep learning models can also enhance the trust and transparency in these systems. Collaboration among machine learning experts, roboticists, and vision scientists can drive the development of innovative algorithms that leverage deep learning to enhance the reliability and accuracy of SLAM-based navigation aids for the visually impaired. This collaboration is essential to bring together the diverse expertise needed to develop effective and practical solutions.
Indoor and outdoor navigation integration Seamless transitions between indoor and outdoor environments are crucial for enhancing the independence and mobility of BVI individuals. However, most studies in our SLR have focused primarily on indoor environments. This limitation arises because indoor environments are often more structured and predictable than outdoor environments, making them easier to map and navigate using SLAM. Outdoor environments, on the other hand, present challenges such as varying lighting conditions, weather changes, and a wider range of obstacles.
Future research should aim to develop solutions that provide unified and consistent navigation experience in both indoor and outdoor settings. This is important because visually impaired individuals need to be able to navigate seamlessly between different environments in their daily lives. A navigation system that only works indoors or outdoors would be of limited use. Researchers should explore the integration of robust sensor fusion techniques and adaptive algorithms capable of handling the different conditions and challenges of these environments. Sensor fusion can combine data from multiple sensors, such as cameras, LiDAR, and IMUs, to provide a more comprehensive and accurate understanding of the environment. Adaptive algorithms can adjust the SLAM system’s parameters in real time to account for changes in lighting, weather, and other environmental factors.
Obstacle detection Achieving detailed knowledge of obstacles and their characteristics is essential for BVI people. Although some studies included in the SLR addressed obstacle detection, the depth and accuracy of the obstacle information provided may still be limited. This limitation stems from the fact that traditional obstacle detection methods often focus on identifying the presence and location of obstacles but may not provide detailed information about their shape, size, or material, which is crucial for visually impaired individuals to make informed decisions during navigation.
@@
To address the need for more detailed and accurate obstacle information, future research should focus on advancing SLAM algorithms to deliver context-aware obstacle detection. This involves integrating semantic understanding with precise spatial measurements, allowing the system to identify and interpret the nature and significance of obstacles accurately. By incorporating semantic understanding, SLAM systems can differentiate between different types of obstacles, such as curbs, stairs, or low-hanging branches, and provide more relevant and actionable information to the user. Precise spatial measurements are essential for accurately estimating the distance, size, and shape of obstacles, enabling visually impaired individuals to navigate safely around them.
@@
In addition, it is crucial to develop algorithms that can learn and adapt to various obstacle types and scenarios. Drawing inspiration from the approaches used in robotics and autonomous drones, such as the real-time metric-semantic SLAM demonstrated by [118], can provide valuable insights. These approaches have demonstrated the feasibility and effectiveness of integrating semantic understanding with precise spatial measurements in real-time SLAM systems. Therefore, future research should prioritize improving both the depth and accuracy of obstacle information, while ensuring robust real-time performance and adaptability to various real-world conditions.
@@
Semantic information integration Integrating semantic information into SLAM algorithms can significantly enhance the performance and robustness of navigation systems for BVI individuals. This information can be used to refine the mapping and localization processes and enhance the overall reliability of navigation in complex environments. For instance, semantic information aids in rejecting outliers during loop-closure detection, which is a crucial SLAM step that identifies and matches previously visited locations. To advance this area, future research should focus on developing advanced techniques for semantic data extraction and integration within the SLAM frameworks. This is necessary because current methods for semantic data extraction and integration may not be efficient enough for real-time SLAM in complex environments. By developing more advanced techniques, researchers can improve the quality and reliability of semantic information used in SLAM, leading to better navigation performance.
Researchers should also explore methods to ensure real-time performance while maintaining the accuracy and detail of the semantic information. Real-time performance is crucial for providing visually impaired users with timely and relevant feedback during navigation. However, processing and integrating semantic information is computationally expensive. Therefore, it is important to develop methods that can balance real-time performance with the accuracy and detail of semantic information.
Realistic dataset creation Another crucial area for future research is the development of realistic and comprehensive datasets tailored for BVI navigation. Although some datasets exist for various SLAM applications, they often do not capture the unique challenges faced by the visually impaired, such as the need to navigate crowded spaces and avoid obstacles at different heights. Future research should focus on creating large-scale, diverse datasets that include various indoor and outdoor settings, different lighting conditions, and dynamic elements. This is important because the lack of such datasets hinders the development and evaluation of SLAM algorithms that are specifically designed for BVI navigation. By creating datasets that reflect real-world challenges faced by visually impaired individuals, researchers can develop more effective and reliable navigation solutions.
Computing resources and battery life The development and deployment of SLAM-based navigation systems for the visually impaired face significant challenges related to computing resources and battery life. SLAM algorithms, particularly when integrated with deep learning models, often demand substantial computing power that can quickly drain battery life and generate heat in mobile devices. This is a critical issue because visually impaired individuals need portable and comfortable navigation aids that can operate for extended periods without overheating or frequent recharging.
Additionally, intensive computations and continuous sensor usage drain the battery life quickly, limiting the usability of the system in real-world daily scenarios. This limitation can significantly hinder the adoption and effectiveness of the SLAM-based navigation systems. Future research should focus on optimizing SLAM algorithms and deep learning models for low-power devices without compromising the accuracy or real-time performance. This is essential for developing energy-efficient solutions that can operate on portable devices with limited battery capacity, ensuring that visually impaired individuals can rely on these systems for extended periods without interruption. Exploring edge computing solutions, developing more efficient neural network architectures, and enhancing battery management techniques could help to address these challenges. These approaches can collectively contribute to reducing the computational burden and power consumption of SLAM-based navigation systems, making them more practical and sustainable for real-world use.
@@
Product development and collaboration Notably, all reviewed approaches were prototypes in the early stages of research and are not yet practical. This might be due to the absence of a unified community or group dedicated to solving the BVI navigation challenges. Much of the work in this domain has been conducted by academic groups or small companies that often fail to produce feasible final products. This underscores a significant future opportunity to develop collaboration and to bridge the gap between research and practical implementation.
Additionally, efforts should be made to develop standardized evaluation metrics and protocols to ensure that the developed systems meet real-world needs and can be effectively transitioned from prototypes to market-ready solutions. Standardized evaluation metrics and protocols are essential to ensure that assistive navigation systems are evaluated consistently and objectively. This can help identify the strengths and weaknesses of different approaches and guide the development of more effective solutions. Encouraging partnerships with technology companies can also accelerate the commercialization process. These partnerships provide the necessary support to bring innovative solutions to the market.
@@ V Conclusion
We investigated the challenging scenarios encountered by SLAM-based navigation systems, which have been addressed in the literature. Additionally, we discussed practical challenges and considerations that affect the usability and adoption of these systems. Furthermore, we analyzed how the proposed SLAM-based solutions improve the mobility and navigation of visually impaired individuals. We evaluated the effectiveness of these solutions in real-world scenarios and assessed the user satisfaction to understand their practical impact on BVI mobility. Finally, we identified gaps, opportunities, and areas of interest that could be explored further in future research, such as addressing challenges in crowded environments, improving real-world applicability, integrating deep learning, and ensuring long-term navigation effectiveness in SLAM-based solutions for visually impaired navigation.